# Turbulence-induced anti-Stokes flow: experiments and theory

### v1

@@ Abstract
We report experimental evidence of an Eulerian-mean flow, $\overline{u}(z)$, created by the interaction of surface waves and tailored ambient sub-surface turbulence, which partly cancels the Stokes drift, $u_{s}(z)$, and present supporting theory. Water-side turbulent velocity fields and Eulerian-mean flows were measured with particle image velocimetry before vs after the passage of a wave group, and with vs without the presence of regular waves. We compare different wavelengths, steepnesses and turbulent intensities. In all cases, a significant change in the Eulerian-mean current is observed, strongly focused near the surface, where it opposes the Stokes drift. The observations support the picture that when waves encounter ambient sub-surface turbulence, the flow undergoes a transition during which Eulerian-mean momentum is redistributed vertically (without changing the depth-integrated mass transport) until a new equilibrium state is reached, wherein the near-surface ratio between $\mathrm{d}\overline{u}/\mathrm{d}z$ and $|\mathrm{d}u_{s}/\mathrm{d}z|$ approximately equals the ratio between the streamwise and vertical Reynolds normal stresses. This accords with a simple statistical theory derived here and holds regardless of the absolute turbulence level, whereas stronger turbulence means faster growth of the Eulerian-mean current. We present a model based on Rapid Distortion Theory which describes the generation of the Eulerian-mean flow as a consequence of the action of the Stokes drift on the background turbulence. Predictions are in qualitative, and reasonable quantitative, agreement with experiments on wave groups, where equilibrium has not yet been reached. Our results could have substantial consequences for predicting the transport of water-borne material in the oceans.
@@ 1 Introduction
Floating and suspended matter of small enough size, such as microplastics (van Sebille, 2020), oil spills (Boufadel et al., 2021), plankton (Hernandez-Carrasco et al., 2018), larvae and nutrients (Röhrs et al., 2014), are transported in the oceans with the Lagrangian-mean current $u_{L}$, equal to the Eulerian-mean current, ${\boldsymbol{u}}(z)$, plus the Stokes drift ${\boldsymbol{u}}_{s}$, i.e., ${\boldsymbol{u}}_{L}(z)={\boldsymbol{u}}(z)+{\boldsymbol{u}}_{s}(z)$. The importance of correctly modelling ocean transport makes prediction of the Lagrangian current a pivotal, but still open, question. For a monochromatic (nearly) linear wave of wavenumber $k_{0}$ and constant amplitude $a$ the Stokes drift velocity along the direction of wave propagation is
$u_{s}(z)=\epsilon_{0}^{2}c(k_{0})\mathrm{e}^{2k_{0}z},$ (1)
@@
where the intrinsic phase velocity (i.e., in the reference system where the surface is at rest) is $c$, the steepness is defined as $\epsilon_{0}=k_{0}a$, and the water is assumed to be deep. The phase velocity and group velocity, $c_{g}$, are then given by
$c(k_{0})=\sqrt{g/k_{0}}\text{ and }c_{g}(k_{0})={\textstyle\frac{1}{2}}\sqrt{g/k_{0}}.$ (2)
@@
A contribution to oceanic transport velocities of magnitude $u_{s}$ can be highly significant and must be taken into account when predicting, e.g., the development of oil spills or the fate of microplastic particles (Hackett et al., 2006; Onink et al., 2019; Cunningham et al., 2022).
@@
The naïve approach for the modeller is to obtain the Lagrangian-mean current by adding the Stokes drift calculated from wave spectra (e.g. Webb & Fox-Kemper, 2011) to the Eulerian flow, which is assumed to be unaffected by waves. However, several field studies have observed that waves appear to have no discernible effect on the Lagrangian-mean current, contrary to theory (Smith, 2006; Lentz et al., 2008). Smith (2006) found that even short wave groups experience an Eulerian current which acts to entirely cancel the Stokes drift at the surface, and that the counter-current is strongly correlated with the presence of Stokes drif, appearing only when the wave group is present and disappears once it has passed. Other observations find that the inclusion of Stokes drift does improve results, however, e.g. Röhrs et al. (2012), who used drifters in coastal waters and employ a coupled wave-ocean model.
@@
Experiments of waves propagating on currents have also yielded results which are inconsistent with a simple addition of Stokes drift. In a careful laboratory study, Monismith et al. (2007) found no change in Lagrangian-mean flow when waves were added, i.e., the Stokes drift is locally cancelled by an equal and opposite Eulerian flow. Moreover, reanalysis of previous experiments by Swan (1990), Jiang & Street (1991), and Thais (1994) supported the same conclusion ( the observation seems to have been made by the same group in the PhD work of Cowen, 1996). These results have remained something of a puzzle, as “[n]o existing theory of wave-current interactions explains this behaviour” as Monismith et al. (2007) put it. Here, we demonstrate a mechanism that could resolve this conundrum at least in part. While being careful not to draw definite conclusions concerning the above-mentioned results, one might remark that some level of turbulence was present alongside the waves in all these experiments: The flow of Swan (1990) passed through a honeycomb flow straightener which will have generated significant turbulence levels, and the mean shear of the flow would provide further production, Jiang & Street (1991) discuss wave-turbulence interactions in their experiments in detail (further detailed by Cheung & Street, 1988), as do Thais & Magnaudet (1996) for the experiment of Thais (1994).
@@
Pearson (2018) predicted with a simple theoretical argument that the interaction between waves and ambient turbulence would, on average, produce an Eulerian-mean current which opposes the Stokes drift near the surface. His paper has received little attention up until now, but in our theoretical work later, we shall draw heavily on his work and apply it to our own settings. Pearson’s prediction shows that the Eulerian-mean current, $\bar{u}(z)$, is opposite and similar to $u_{s}(z)$ near the surface but changes sign beneath the wave-influenced surface layer and integrates to zero as a function of depth. Although the wave-turbulence-induced Eulerian-mean current incurs no net change in mass transport, it will partly cancel the Stokes drift at the surface, and we therefore refer to it as an ‘anti-Stokes’ current.
@@
The picture which emerges is that when waves and turbulence first meet, the combined flow goes through a transient ‘spin-up’ period before a new quasi-equilibrium is reached, which includes the Eulerian anti-Stokes current. We review and extend statistical theory for both the transient and steady stages, and derive a theory based on Rapid Distortion Theory (RDT) which captures the underlying physics of the ‘spin-up’ period and shows that the Eulerian acceleration of the current will exhibit approximately the same depth-dependent behaviour as the resulting current that we observe. The process is intimately related to the so-called ‘CL2’ mechanism which creates Langmuir circulation (Craik & Leibovich, 1976). It is worth noting that the CL2 mechanism, as lucidly reviewed by Leibovich (1983), requires the presence of an Eulerian-mean flow with slope (i.e., $\mathrm{d}\bar{u}/\mathrm{d}z$) of the same sign as that of the Stokes drift profile, whereas the anti-Stokes current induced by wave-turbulence interaction, $\bar{u}(z)$, has opposite slope. Thus, $(\mathrm{d}\bar{u}/\mathrm{d}z)\cdot(\mathrm{d}u_{s}/\mathrm{d}z)<0$, which implies that the induced current tends to stabilise the combined system with respect to the CL2 instability.
@@
In fact, there have been indications for several decades that the same physical phenomenon has been at play in wave-current experiments performed in the context of studying the bottom boundary layer in shallow wave-current flows motivated by understanding sediment transport (thus not highlighting the significance for ocean modelling). A string of independent measurements of the mean Eulerian flow in the presence and absence of waves by van Hoften & Karaki (1977); Bakker & van Doorn (1978); van Doorn (1981); Kemp & Simons (1982, 1983); Rashidi et al. (1992); Klopman (1994); Mathisen & Madsen (1996) and Singh & Debnath (2017), all showed that the waves caused a significant alteration of the mean flow near the surface, adding a contribution in the direction opposite to wave propagation in the near-surface region. More recent experiments also report the same (Zhang & Simons, 2019; Peruzzi et al., 2021). In addition to making the same observation, Umeyama (2005, 2009) found in his experiments that the vertical structure of the streamwise-vertical Reynolds shear stress depends strongly on the wave-propagation direction in the near-surface region.
@@
Since these studies considered shallow currents where waves affect the bottom boundary layer, a direct comparison with our experiment in deep water is dubious, yet subsequent theoretical analysis gives reason to suspect a close connection. Nielsen & You (1996) and Dingemans et al. (1996) propose two early explanations for the difference in Eulerian-mean current; the former relies on a force balance on average including the mean stress from waves and turbulence represented by eddy viscosity, the latter on the creation of streamwise rolls due to the Craik–Leibovich vortex force due to the sidewall boundary layers, whose presence was observed by Klopman (1994). Groeneweg & Klopman (1998) developed a more sophisticated theory based on Generalised Lagrangian-Mean (GLM) theory, similar in spirit to the physical process we consider herein, but their analysis is not easy to compare with ours since it involves a set of coupled nonlinear differential equations and a complex turbulence model. Interestingly, Groeneweg & Battjes (2003) reconcile all three descriptions, at least qualitatively, by extending their GLM theory to three dimensions. Huang & Mei (2003) provided a careful theory, once more primarily concerned with the effect of waves on the bottom boundary layer, but also shedding light on the observed changes near the surface. They, too, use the simple eddy viscosity model of turbulence. In particular, they conclude that the mean wave-induced shear stress near the free surface is opposite in direction to the wave propagation, and is “due largely to the distortion of eddy viscosity near the surface”. With a simple mixing-length model, Umeyama (2005, 2009) as well as Yang et al. (2006) find reasonable agreement with experiments. Olabarrieta et al. (2010) devise a simplified numerical model to avoid the restriction of low-steepness waves, and the perturbation theory of Tambroni et al. (2015) yields a model able to predict the Eulerian-mean velocity profile throughout the water column; both of the latter employ depth-dependent eddy-viscosity models. Crucially, all of these many model explanations depend on the simultaneous presence of waves and turbulence to explain the change in mean flow.
@@
When reviewing previous numerical studies of waves in the presence of turbulence, one can also find evidence of the same Eulerian-mean current creation that we observe experimentally, even though the authors themselves have not discussed its significance especially. Borue et al. (1995) merely remark that the change in mean current near the surface should be studied further, while Kawamura (2000) notes that the changes in current are higher the larger the Stokes drift magnitude, and discusses consequences for (Langmuir) turbulence production. Most strikingly, Fujiwara et al. (2020) find Eulerian-mean velocity profiles under Langmuir turbulence which are qualitatively very similar to those we measure under regular waves (their figures 6a and 9a) and report in section 2.1.2, but do not discuss this point particularly.
@@
Eulerian-mean flow driven by waves also occurs without pre-existing turbulence or vorticity from two separate mechanisms. First, to compensate for the divergence of Stokes drift on the group scale, Stokes drift also in otherwise quiescent water must be accompanied by an Eulerian return flow (e.g., Longuet-Higgins & Stewart, 1962; van den Bremer & Taylor, 2016) for wave groups. In deep water, the depth-integrated Stokes drift and the depth-integrated Eulerian return flow are equal and opposite, and mass is preserved globally, not locally. This phenomenon is reviewed further in section 4.1. The Stokes drift profile is highly concentrated near the surface whereas the return flow varies slowly with depth. The Eulerian-mean “anti-Stokes” current we observe also varies rapidly with depth and cannot be explained by this mass conservation mechanism. Second, there is also surface streaming driven by viscosity confined to a thin viscous boundary layer beneath the surface, resulting from the imparting of wave momentum to the fluid as the waves decay (Longuet-Higgins, 1953; Craik, 1982). Tsai et al. (2017) and, recently, Fujiwara (2024) studied how this current can also interact with waves to generate small-scale turbulence of Langmuir type via the CL2 mechanism. The viscous sub-surface layer is likely thinner than what our measurements can resolve, and, additionally, the resulting Eulerian-mean flow is directed along, not against, the direction of wave propagation. Third, the Earth’s rotation causes a wave-induced Eulerian-mean flow that can exactly cancel the Stokes drift (for periodic waves and in the absence of viscosity), which is also known as the anti-Stokes flow (Hasselmann, 1970). While this rotation-induced anti-Stokes flow could explain field observations (Lentz et al., 2008; Röhrs et al., 2012), it cannot explain experimental results discussed above and presented herein as our characteristic timescales are vastly smaller than Earth’s period of rotation (the Rossby number is large). Hence, neither of these three mechanisms can explain the observations just mentioned, nor those by, e.g., Monismith et al. (2007) or indeed those we report herein.

### v2

@@ Abstract
We report experimental evidence of an Eulerian-mean flow, $\overline{u}(z)$, created by the interaction of surface waves and tailored ambient sub-surface turbulence, which partly cancels the Stokes drift, $u_{s}(z)$, and present supporting theory. Water-side turbulent velocity fields and Eulerian-mean flows were measured with particle image velocimetry before vs after the passage of a wave group, and with vs without the presence of regular waves. We compare different wavelengths, steepnesses and turbulent intensities. In all cases, a significant change in the Eulerian-mean current is observed, strongly focused near the surface, where it opposes the Stokes drift. The observations support the picture that when waves encounter ambient sub-surface turbulence, the flow undergoes a transition during which Eulerian-mean momentum is redistributed vertically (without changing the depth-integrated mass transport) until a new equilibrium state is reached, wherein the near-surface ratio between $|\mathrm{d}\overline{u}/\mathrm{d}z|$ and $|\mathrm{d}u_{s}/\mathrm{d}z|$ approximately equals the ratio between the streamwise and vertical Reynolds normal stresses. This accords with a simple statistical theory derived here and holds regardless of the absolute turbulence level, whereas stronger turbulence means faster growth of the Eulerian-mean current. We present a model based on Rapid Distortion Theory which describes the generation of the Eulerian-mean flow as a consequence of the action of the Stokes drift on the background turbulence. Predictions are in qualitative, and reasonable quantitative, agreement with experiments on wave groups, where equilibrium has not yet been reached. Our results could have substantial consequences for predicting the transport of water-borne material in the oceans.
@@ 1 Introduction
Floating and suspended matter of small enough size, such as microplastics (van Sebille, 2020), oil spills (Boufadel et al., 2021), plankton (Hernandez-Carrasco et al., 2018), larvae and nutrients (Röhrs et al., 2014), are transported in the oceans with the Lagrangian-mean current $\bar{u}_{L}$, equal to the Eulerian-mean current, $\bar{{\boldsymbol{u}}}(z)$, plus the Stokes drift ${\boldsymbol{u}}_{s}$, i.e., $\bar{{\boldsymbol{u}}}_{L}(z)=\bar{{\boldsymbol{u}}}(z)+{\boldsymbol{u}}_{s}(z)$. The importance of correctly modelling ocean transport makes prediction of the Lagrangian current a pivotal, but still open, question. For a monochromatic (nearly) linear wave of wavenumber $k_{0}$ and constant amplitude $a$ propagating in deep water the Stokes drift velocity along the direction of wave propagation is
$u_{s}(z)=(ak_{0})^{2}c(k_{0})\mathrm{e}^{2k_{0}z},$ (1)
@@
where the intrinsic phase velocity (i.e., in the reference system where the surface is at rest) is $c$. The phase velocity and group velocity, $c_{g}$, are given by
$c(k_{0})=\sqrt{g/k_{0}}\text{ and }c_{g}(k_{0})={\textstyle\frac{1}{2}}\sqrt{g/k_{0}}$ (2)
@@
where $g$ is the gravitational acceleration. A contribution to oceanic transport velocities of magnitude $u_{s}$ can be highly significant and must be taken into account when predicting, e.g., the development of oil spills or the fate of microplastic particles (Hackett et al., 2006; Onink et al., 2019; Cunningham et al., 2022).
@@
The naïve approach for the modeller is to obtain the Lagrangian-mean current by adding the Stokes drift calculated from wave spectra to the Eulerian flow, which is assumed to be unaffected by waves. However, several field studies have observed that waves appear to have no discernible effect on the Lagrangian-mean current, contrary to theory (Smith, 2006; Lentz et al., 2008). Smith (2006) found that even short wave groups experience an Eulerian-mean current, which acts to entirely cancel the Stokes drift at the surface, and that the counter-current is strongly correlated with the presence of Stokes drift – it appears only when the wave group is present and disappears once it has passed. Other observations find that the inclusion of Stokes drift does improve results, however, e.g. Röhrs et al. (2012), who used drifters in coastal waters and employ a coupled wave–ocean model.
@@
Experiments of waves propagating on currents have also yielded results which are inconsistent with a simple addition of Stokes drift. In a careful laboratory study, Monismith et al. (2007) found no change in Lagrangian-mean flow when waves were added, i.e., the Stokes drift is locally cancelled by an equal and opposite Eulerian flow. Moreover, reanalysis of previous experiments by Swan (1990b, a), Jiang & Street (1991), and Thais (1994) supported the same conclusion (the observation seems to have been made also in the PhD work of Cowen, 1996). These results have remained something of a puzzle, as “[n]o existing theory of wave–current interactions explains this behaviour” as Monismith et al. (2007) put it. Here, we demonstrate a mechanism that could resolve this conundrum at least in part.
While being careful not to draw definite conclusions concerning the above-mentioned results, one might remark that some level of turbulence was likely to have been present alongside the waves in all these experiments. The flow of Swan (1990b) passed through a honeycomb flow straightener which will have generated significant turbulence levels, and the mean shear of the flow would provide further turbulence production, Jiang & Street (1991) discuss wave–turbulence interactions in their experiments in detail (further detailed by Cheung & Street, 1988), as do Thais & Magnaudet (1996) for the experiment of Thais (1994). The experiment of Swan (1990a) did not have imposed wind or current. Since measurements were made after the wavemaker had run for many hours, one may speculate that turbulence might well have been present, created, e.g., in the oscillating bottom boundary layer (the depth was less than a third of a wavelength) or thermal convection, with which the waves would have had ample time to interact, the flume being $18$ m long. Naturally, this can be no more than speculation so long afterwards.
@@
Pearson (2018) predicted with a theoretical argument that the interaction between waves and ambient turbulence would, on average, produce an Eulerian-mean current which opposes the Stokes drift near the surface. His paper has received little attention up until now, but in our theoretical work later, we shall draw heavily on his work and apply it to our own settings. Pearson’s prediction shows that the Eulerian-mean current, $\bar{u}(z)$, is opposite and similar to $u_{s}(z)$ near the surface but changes sign beneath the wave-influenced surface layer and integrates to zero as a function of depth. Although the wave–turbulence-induced Eulerian-mean current incurs no net change in mass transport, it will partly cancel the Stokes drift at the surface, and we therefore refer to it as an ‘anti-Stokes’ current.
The picture which emerges is that, when waves and turbulence first meet, the combined flow goes through a transient ‘spin-up’ period before a new quasi-equilibrium is reached, which includes the Eulerian anti-Stokes current. In our set-up, turbulence is created at the inlet of the mean flow and as it moves downstream towards the point of measurement it encounters waves and develops together with the velocity profile. When the time of effective interaction is short, the result of partial spin-up is captured, while turbulence moving in the presence of waves for sufficiently long could reach a quasi-steady state before it is measured. The conditions for ‘short’ and ‘long’ interaction time are satisfied in our experiments, fo wave groups (Experiment 1) and regular waves (Experiments 2 and 3), respectively.
@@
We review and extend statistical theory for both the transient and steady stages, and derive a theory based on Rapid Distortion Theory (RDT) which captures the underlying physics of the ‘spin-up’ period and shows that the Eulerian acceleration of the current will exhibit approximately the same depth-dependent behaviour as the resulting current that we observe. The process is intimately related to the so-called ‘CL2’ mechanism which creates Langmuir circulation (Craik & Leibovich, 1976). It is worth noting that the CL2 mechanism, as lucidly reviewed by Leibovich (1983), requires the presence of an Eulerian-mean flow with slope (i.e., $\mathrm{d}\bar{u}/\mathrm{d}z$) of the same sign as that of the Stokes drift profile, whereas the anti-Stokes current induced by wave–turbulence interaction, $\bar{u}(z)$, has opposite slope. Thus, $(\mathrm{d}\bar{u}/\mathrm{d}z)\cdot(\mathrm{d}u_{s}/\mathrm{d}z)<0$, which implies that the induced current tends to stabilise the combined system with respect to the CL2 instability.
1.1 Possible previous observations
@@
Indications are that the same physical phenomenon has been at play in wave–current experiments performed in the context of studying the bottom boundary layer in shallow wave–current flows, motivated by understanding sediment transport (thus not highlighting the significance for ocean modelling). A string of independent measurements of the mean Eulerian flow in the presence and absence of waves by van Hoften & Karaki (1977); Bakker & van Doorn (1978); van Doorn (1981); Kemp & Simons (1982, 1983); Rashidi et al. (1992); Klopman (1994); Mathisen & Madsen (1996) and Singh & Debnath (2017), all showed that the waves caused a significant alteration of the mean flow near the surface, adding a contribution in the direction opposite to wave propagation in the near-surface region. More recent experiments also report the same (Zhang & Simons, 2019; Peruzzi et al., 2021). In addition to making the same observation, Umeyama (2005, 2009) found in his experiments that the vertical structure of the streamwise–vertical Reynolds shear stress depends strongly on the wave-propagation direction in the near-surface region.
@@
Since these studies considered shallow currents where waves affect the bottom boundary layer, a direct comparison with our experiment in deep water is dubious, yet subsequent theoretical analysis gives reason to suspect a close connection. Nielsen & You (1996) and Dingemans et al. (1996) propose two early explanations for the difference in Eulerian-mean current; the former relies on a force balance on average including the mean stress from waves and turbulence represented by eddy-viscosity, the latter on the creation of streamwise rolls due to the Craik–Leibovich vortex force due to the sidewall boundary layers, whose presence was observed by Klopman (1994). Groeneweg & Klopman (1998) developed a more sophisticated theory based on Generalised Lagrangian-Mean (GLM) theory, similar in spirit to the physical process we consider herein, but their analysis is not easy to compare with ours since it involves a set of coupled nonlinear differential equations and a complex turbulence model. Interestingly, Groeneweg & Battjes (2003) reconcile all three descriptions, at least qualitatively, by extending their GLM theory to three dimensions. Huang & Mei (2003) provided a careful theory, once more primarily concerned with the effect of waves on the bottom boundary layer, but also shedding light on the observed changes near the surface. They, too, use the simple eddy viscosity model of turbulence. In particular, they conclude that the mean wave-induced shear stress near the free surface is opposite in direction to the wave propagation, and is “due largely to the distortion of eddy viscosity near the surface”. With a simple mixing-length model, Umeyama (2005, 2009) as well as Yang et al. (2006) find reasonable agreement with experiments. Olabarrieta et al. (2010) devise a simplified numerical model to avoid the restriction of low-steepness waves, and the perturbation theory of Tambroni et al. (2015) yields a model able to predict the Eulerian-mean velocity profile throughout the water column; both of the latter employ depth-dependent eddy-viscosity models. Crucially, all of these many model explanations depend on the simultaneous presence of waves and turbulence to explain the change in mean flow.
@@
When reviewing previous numerical studies of waves in the presence of turbulence, one can also find evidence of the same Eulerian-mean current creation that we observe experimentally, even though the authors themselves have not discussed its significance especially. Borue et al. (1995) merely remark that the change in mean current near the surface should be studied further, while Kawamura (2000) notes that the changes in current are higher the larger the Stokes drift magnitude, and discusses consequences for (Langmuir) turbulence production. Also Fujiwara et al. (2020) find Eulerian-mean velocity profiles under Langmuir turbulence which are qualitatively very similar to those we measure under regular waves (their figures 6a and 9a) and report in section 2.1.2, but do not discuss this point particularly.
1.2 Other wave-driven currents
@@
Eulerian-mean flow driven by waves also occurs without pre-existing turbulence or vorticity from two separate mechanisms. First, to compensate for the divergence of Stokes drift on the group scale, Stokes drift in otherwise quiescent water must also be accompanied by an Eulerian return flow (e.g., Longuet-Higgins & Stewart, 1962; van den Bremer & Taylor, 2016) for wave groups. In deep water, the depth-integrated Stokes drift and the depth-integrated Eulerian return flow are equal and opposite, and mass is preserved globally, not locally. This phenomenon is reviewed further in section 4.1. The Stokes drift profile is highly concentrated near the surface whereas the return flow varies slowly with depth. The Eulerian-mean “anti-Stokes” current we observe varies rapidly with depth and cannot be explained by this mass conservation mechanism. Second, there is also surface streaming driven by viscosity confined to a thin viscous boundary layer beneath the surface, resulting from the imparting of wave momentum to the fluid as the waves decay (Longuet-Higgins, 1953; Craik, 1982). Tsai et al. (2017) and, recently, Fujiwara (2024) studied how this current can also interact with waves to generate small-scale turbulence of Langmuir type via the CL2 mechanism. The viscous sub-surface layer is likely thinner than what our measurements can resolve, and, additionally, the resulting Eulerian-mean flow is directed along, not against, the direction of wave propagation. Third, the Earth’s rotation causes a wave-induced Eulerian-mean flow that can exactly cancel the Stokes drift (for periodic waves and in the absence of viscosity), which is also known as the anti-Stokes flow (Hasselmann, 1970). While this rotation-induced anti-Stokes flow could explain some field observations (Lentz et al., 2008; Röhrs et al., 2012), it cannot explain experimental results discussed above and presented herein as our characteristic timescales are vastly smaller than Earth’s period of rotation (the Rossby number is large). Hence, neither of these three mechanisms can explain the observations just mentioned, nor those by, e.g., Monismith et al. (2007) or indeed those we report herein.